OpenStudy (anonymous):

Can anyone help? Plans for a shed call for a rectangular floor with a perimeter of 252 ft. Thel ength is two times the width. What is the length and the width of the floor?

OpenStudy (anonymous):

P = 2l + 2w./252 = 2(2x) + 2(x)/252 = 4x + 2x/252 = 6x/x = 42 2x = 84. So length is 84 ft and the width is 42.

OpenStudy (anonymous):

Did that help?

OpenStudy (anonymous):

Got booted off - sorry. Thank you very much!

OpenStudy (anonymous):

I posted another one -can u help me? Math not my strongsuit and I am sick on top of it. "o(